# Relic gravitational waves from primordial gravitational collapses

### v1

@@ Abstract
It has long been expected Hawke and Stewart (2002); Musco et al. (2005); Musco and Miller (2013); Bloomfield et al. (2015); García-Bellido et al. (2021); Carr et al. (2021) that an outgoing sound shell Deng and Vilenkin (2017); Deng et al. (2018); Deng (2020a, b) will form along with primordial black hole (PBH) formation from gravitational collapse of a sufficiently large density perturbation of horizon size. In this Letter, we will report a new source of SGWBs induced by the superposition of sound waves from gravitational collapses of large density perturbations of horizon size in the early Universe via full numerical relativity. The details of numerical simulations are presented separately in an accompanying long paper Ning et al. on the long-time evolution of such large density perturbations. This Letter will focus on the GW phenomenology even if these density perturbations are not large enough to form PBHs.
Sound waves from density perturbations. PBHs can form in the early Universe if the perturbation amplitude exceeds a critical threshold Carr and Hawking (1974); Hawking (1971). For those density perturbations of horizon size that are not sufficiently dense to result in PBH formation, bounces occur, and a shell of density contrast dubbed sound wave propagates outwards, as shown in previous numerical simulation results Hawke and Stewart (2002); Musco et al. (2005); Bloomfield et al. (2015); Musco and Miller (2013). In the accompanying long paper Ning et al. , we find that the sound wave could also form for the case with PBH formation, and hence, we will follow the numerical treatment of the PBH formation.
@@
where $r_{m}$ is the length scale of the perturbation and $\mathcal{A}$ denotes its peak amplitude. The curvature perturbation can induce density contrast $\delta_{m}\equiv\delta(r_{m},t_{m})$ at the horizon crossing time $t_{m}$ via $\delta_{m}=\frac{3(1+\omega)}{e(5+3\omega)}\mathcal{A}r_{m}^{2}$ Musco (2019). With the above initial profile, the threshold for PBH formation is $\delta_{c}\simeq 0.4977$ in the radiation-dominated (RD) era with $\omega=1/3$ as obtained in our accompanying long paper Ning et al. , which numerically evolves the density perturbation collapse for the first time in the long-time limit. We modify the code developed in Ref. Escrivà (2020) and perform all of our simulations in the RD era, but the generalization to other $\omega$ should be straightforward.
@@
Observational perspectives. As shown in the Supplementary Material, the peak frequency in the GW spectrum is about a quarter Ning et al. of comoving thickness $d_{c}$ of underdense shell at collisions so that we can define a nearly constant $m\equiv d_{c}f_{\mathrm{peak}}\simeq\mathcal{O}(1)$. On the other hand, the sound shell model has found that the peak frequency is roughly the sound-shell thickness at free propagation, which has been found in our accompanying long paper Ning et al. that $d_{c}$ will approach a constant proportionality to the comoving Hubble horizon $\mathcal{H}_{f}$ at reentry of initial superhorizon perturbations. Therefore, after characterizing $d_{c}\equiv n\mathcal{H}_{f}^{-1}$ with $n\simeq\mathcal{O}(1)$, we can estimate the present-day peak frequency $f_{0}=(a_{f}/a_{0})f_{\mathrm{peak}}=(a_{f}/a_{0})(m/n)(H_{f}/2\pi)$ in terms of PBH mass $M_{\mathrm{PBH}}=4\pi\gamma M_{\mathrm{Pl}}^{2}/H_{f}$ Carr and Hawking (1974); Hawking (1971); Sasaki et al. (2018) as
@@
Constraints on PBH abundance. One application of this new source of SGWBs is to constrain the abundance of PBHs, especially those that have already evaporated today. To investigate the PBH abundance, it is more convenient to use the density fraction of PBHs at the epoch of their formation, denoted as
@@
Conclusions and discussions. It is of essential importance to thoroughly study all possible GW sources so as to fully understand the SGWBs in GW detectors. In this Letter, we have proposed for the first time a new SGWB sourced by the collisions of sound shells from density perturbation collapse, providing an alternative probe into the evaporated PBH mass and abundance.
@@ Appendix A Abstract
This Supplemental Material is prepared for the technical details of the dimensionless gravitational-wave (GW) power spectrum from the primordial gravitational collapse. The details for implementing long-time numerical simulations of primordial gravitational collapses of density perturbations are prepared in a separate accompanying paper Ning et al. .

### v2

@@ Abstract
It has long been expected Hawke and Stewart (2002); Musco et al. (2005); Musco and Miller (2013); Bloomfield et al. (2015); García-Bellido et al. (2021); Carr et al. (2021) that an outgoing sound shell Deng and Vilenkin (2017); Deng et al. (2018); Deng (2020a, b) will form along with primordial black hole (PBH) formation from gravitational collapse of a sufficiently large density perturbation of horizon size. In this Letter, we will report a new source of SGWBs induced by the superposition of sound waves from gravitational collapses of large density perturbations of horizon size in the early Universe via full numerical relativity. The details of numerical simulations are presented separately in an accompanying long paper Ning et al. (2025) on the long-time evolution of such large density perturbations. This Letter will focus on the GW phenomenology even if these density perturbations are not large enough to form PBHs.
Sound waves from density perturbations. PBHs can form in the early Universe if the perturbation amplitude exceeds a critical threshold Carr and Hawking (1974); Hawking (1971). For those density perturbations of horizon size that are not sufficiently dense to result in PBH formation, bounces occur, and a shell of density contrast dubbed sound wave propagates outwards, as shown in previous numerical simulation results Hawke and Stewart (2002); Musco et al. (2005); Bloomfield et al. (2015); Musco and Miller (2013). In the accompanying long paper Ning et al. (2025), we find that the sound wave could also form for the case with PBH formation, and hence, we will follow the numerical treatment of the PBH formation.
@@
where $r_{m}$ is the length scale of the perturbation and $\mathcal{A}$ denotes its peak amplitude. The curvature perturbation can induce density contrast $\delta_{m}\equiv\delta(r_{m},t_{m})$ at the horizon crossing time $t_{m}$ via $\delta_{m}=\frac{3(1+\omega)}{e(5+3\omega)}\mathcal{A}r_{m}^{2}$ Musco (2019). With the above initial profile, the threshold for PBH formation is $\delta_{c}\simeq 0.4977$ in the radiation-dominated (RD) era with $\omega=1/3$ as obtained in our accompanying long paper Ning et al. (2025), which numerically evolves the density perturbation collapse for the first time in the long-time limit. We modify the code developed in Ref. Escrivà (2020) and perform all of our simulations in the RD era, but the generalization to other $\omega$ should be straightforward.
@@
Observational perspectives. As shown in the Supplementary Material, the peak frequency in the GW spectrum is about a quarter Ning et al. (2025) of comoving thickness $d_{c}$ of underdense shell at collisions so that we can define a nearly constant $m\equiv d_{c}f_{\mathrm{peak}}\simeq\mathcal{O}(1)$. On the other hand, the sound shell model has found that the peak frequency is roughly the sound-shell thickness at free propagation, which has been found in our accompanying long paper Ning et al. (2025) that $d_{c}$ will approach a constant proportionality to the comoving Hubble horizon $\mathcal{H}_{f}$ at reentry of initial superhorizon perturbations. Therefore, after characterizing $d_{c}\equiv n\mathcal{H}_{f}^{-1}$ with $n\simeq\mathcal{O}(1)$, we can estimate the present-day peak frequency $f_{0}=(a_{f}/a_{0})f_{\mathrm{peak}}=(a_{f}/a_{0})(m/n)(H_{f}/2\pi)$ in terms of PBH mass $M_{\mathrm{PBH}}=4\pi\gamma M_{\mathrm{Pl}}^{2}/H_{f}$ Carr and Hawking (1974); Hawking (1971); Sasaki et al. (2018) as
@@
Constraints on PBH abundance. One application of this new source of SGWBs is to constrain the abundance of PBHs, especially those that have already evaporated today (see also Ref. Franciolini and Pani (2023)). To investigate the PBH abundance, it is more convenient to use the density fraction of PBHs at the epoch of their formation,
@@
Conclusions and discussions. It is of essential importance to thoroughly study all possible GW sources so as to fully understand the SGWBs in GW detectors. In this Letter, we have proposed a new SGWB sourced by the collisions of sound shells from density perturbation collapse, providing an alternative probe into the evaporated PBH mass and abundance. We note here that a previous work De Luca et al. (2019) has also studied the GW production but during the non-spherical collapses of sub-critical scalar perturbations, which is overwhelmed by those scalar-induced GWs (SIGW) at second-order in perturbations Ananda et al. (2007); Baumann et al. (2007) (see Ref. Domènech (2021) for a recent review and references therein). We stress here that our newly proposed SGWB is complementary to both of the above two GW sources, as the SIGW is mainly produced at horizon crossing, while our SGWB is generated long after horizon crossing when sound waves from gravitational collapses are in full superpositions.
@@ Appendix A Abstract
This Supplemental Material is prepared for the technical details of the dimensionless gravitational-wave (GW) power spectrum from the primordial gravitational collapse and its comparison to the scalar-induced GWs at second-order in perturbations. The details for implementing long-time numerical simulations of primordial gravitational collapses of density perturbations are prepared in a separate accompanying paper Ning et al. (2025).
@@ Appendix D Appendix C: The peak behavior of the GW power spectrum
Appendix E Appendix D: Comparison with Scalar Induced Gravitational Waves (SIGW)
It is well known that enhanced curvature perturbations can also induce GWs at second-order in perturbations Ananda et al. (2007); Baumann et al. (2007). Therefore, we should compare the amplitude of this scalar-induced GW (SIGW) with our GW generated by sound waves (SWGW). Before diving into the comparison, we should stress that our mechanism applies not only to the case with curvature perturbations but also to other cases that generate sound waves, such as that from Ref. Deng et al. (2018). As a rough estimation, we will use a Gaussian probability distribution function (PDF) for the density perturbations,
$\mathscr{P}_{\mathrm{G}}[\delta]=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{\delta^{2}}{2\sigma^{2}}},$ (52)
where $\sigma^{2}$ is the variance of the distribution calculated by the power spectrum of the density perturbation as
$\sigma^{2}=\langle\delta^{2}\rangle=\int_{0}^{\infty}\mathrm{d}\ln q\mathcal{W}^{2}(q,R)\mathcal{P}_{\delta}(q).$ (53)
Here, $\mathcal{W}$ denotes the Fourier transform of a real space window function and $R$ is the characteristic scale $R\approx\mathcal{H}^{-1}$. In the following calculations, we will choose a top hat window function for a simple illustration,
$\mathcal{W}(q,R)=\frac{3\sin(qR)-3qR\cos(qR)}{(qR)^{3}}.$ (54)
The relation between the power spectrum of density perturbation and curvature perturbation at linear order is given by Josan et al. (2009)
$\mathcal{P}_{\delta}(k)\simeq\frac{16}{81}\left(\frac{k}{aH}\right)^{4}\mathcal{P}_{\mathcal{R}}(k),$ (55)
which, after combined with (53), can be used to relate $\mathcal{P}_{\mathcal{R}}(k)$ with $\sigma^{2}$. Then, using the Press-Schechter model Press and Schechter (1974), one can estimate $\beta$ in terms of $\sigma$ as
$\beta\equiv\int_{\Delta_{c}}^{\infty}\mathscr{P}[\delta]\mathrm{d}\delta.$ (56)
In the following, we will ignore the actual relation between the density contrast $\Delta_{c}$ and $\delta_{c}$, and we will simply choose $\Delta_{c}=\delta_{m}$ for our cases. For a delta-peak curvature perturbation power spectrum $\mathcal{P}_{\mathcal{R}}=A\delta(\ln(k/k_{*}))$, we find the amplitude $A\simeq 10^{-2}$ for all cases, thus, we will choose $A=10^{-2}$ for our rough estimation. Using a more realistic Gaussian power spectrum,
$\mathcal{P}_{\mathcal{R}}(k)=\frac{A(k/k_{*})^{3}}{\sqrt{2\pi}\epsilon}e^{-\frac{((k/k_{*})-1)^{2}}{2\epsilon^{2}}},$ (57)
and choosing $A=10^{-2},\epsilon=10^{-1},k_{*}=\mathcal{H}_{f}=(10r_{H})^{-1}$ for a fair comparison, we can calculate the corresponding SIGW and SWGWs as shown in Fig. 6. As one can see, besides the slightly different peak frequencies, although the amplitude of SIGW is much larger than the corresponding SWGWs at low frequencies, the high-frequency part of SWGWs might be detectable, as there is a high-frequency cutoff in SIGW. We further note here that, although the SIGW seems to be quite similar to our SWGW in nature, as they are all generated from the scalar perturbations in the early Universe, their production eras are different: the SIGW is mainly produced at the time of horizon re-entry, while our SWGW is mainly produced long after horizon re-entry when the sound waves from gravitational collapses are in full superpositions.
Figure 6: The GW energy density spectra of scalar-induced gravitational waves (pink) and our gravitational waves from sound waves of primordial gravitational collapses.